# Huan Zhang

### v1

@@ Abstract
Keywords:
Differential system Inertial algorithm Non-smooth convex optimization Iterates convergence
MSC:
34D0537N40 90C25
@@ 1 Introduction
$\left\{\begin{array}[]{ll}&\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}~~{f(x)}\\ &\mbox{s.t.}~~Ax=b,\end{array}\right$. (3)
@@
$\begin{cases}\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)+\theta t\dot{\lambda}(t)\right)=0,\\ \ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta t\dot{x}(t),\lambda(t)\right)=0,\end{cases}$ (4)
@@
$\begin{cases}\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)+\theta t\dot{\lambda}(t)\right)=0,\\ \ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta t\dot{x}(t),\lambda(t)\right)=0,\end{cases}$ (5)
@@
$\begin{cases}\ddot{x}(t)+r\dot{x}(t)+\beta(t)\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta\dot{x}(t),\lambda(t)\right)=0,\end{cases}$ (6)
@@
$\left\{\begin{array}[]{ll}&\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}~~{f(x)+g(x)}\\ &\mbox{s.t.}~~Ax=b,\end{array}\right$. (9)
@@
$\begin{cases}\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\partial_{x}\mathcal{L}_{\rho}\left(x(t),\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t)\right)\ni 0,\\ \ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\rho}\left(x(t)+\frac{t}{\alpha-1}\dot{x}(t),\lambda(t)\right)=0,\end{cases}$ (10)
@@ 2 Preliminaries
$(x^{*},\lambda^{*})\in\mathcal{S}\Leftrightarrow\begin{cases}0\in\partial_{x}\mathcal{L}_{\rho}(x^{*},\lambda^{*})=\nabla f(x^{*})+\partial g(x^{*})+A^{*}\lambda^{*},\\ 0=\nabla_{\lambda}\mathcal{L}_{\rho}(x^{*},\lambda^{*})=Ax^{*}-b,\end{cases}$
@@ 3 Fast Convergence Rates for Differential System (10)
$\begin{cases}\ddot{x}_{\gamma}(t)+\frac{\alpha}{t}\dot{x}_{\gamma}(t)+\beta(t)\Big{(}\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\\ ~~~~+A^{*}\left(\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right)+\rho A^{*}\left(Ax_{\gamma}(t)-b\right)\Big{)}=0,\\ \ddot{\lambda}_{\gamma}(t)+\frac{\alpha}{t}\dot{\lambda}_{\gamma}(t)-\beta(t)\left(A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b\right)=0.\end{cases}$ (13)
@@
$\begin{array}[]{rl}\mathcal{E}_{\gamma}(t)=&\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &+\frac{1}{2}\lVert x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\rVert^{2}+\frac{1}{2}\lVert\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\rVert^{2}.\end{array}$ (14)
Clearly, $\mathcal{E}_{\gamma}(t)\geq 0$, $\forall t\geq t_{0}$. Then,
$\begin{array}[]{rl}\dot{\mathcal{E}}_{\gamma}(t)&=\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}\left(x_{\gamma}(t),\lambda^{*}\right)-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~~+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\langle\nabla_{x}\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*}),\dot{x}_{\gamma}(t)\rangle\\ &~~~~+\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle\\ &~~~~+\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle\\ &=\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}\left(x_{\gamma}(t),\lambda^{*}\right)-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~~+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right),\dot{x}_{\gamma}(t)\right\rangle\\ &~~~~+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle\lambda^{*},A\dot{x}_{\gamma}(t)\right\rangle+\frac{\rho t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle Ax_{\gamma}(t)-b,A\dot{x}_{\gamma}(t)\right\rangle\\ &~~~~+\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle\\ &~~~~+\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle.\end{array}$
@@
$\begin{array}[]{rl}&\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle\\ &=-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle\\ &~~-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\right.\\ &~~\left.A^{*}\left(\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right)+\rho A^{*}\left(Ax_{\gamma}(t)-b\right)\right\rangle\\ &=-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle\\ &~~-\frac{t\beta(t)}{\alpha-1}\left\langle A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b,\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right\rangle\\ &~~-\frac{\rho t\beta(t)}{\alpha-1}\left\langle A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b,Ax_{\gamma}(t)-b\right\rangle,\end{array}$
@@
$\begin{array}[]{rl}&\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle\\ &=\frac{t\beta(t)}{\alpha-1}\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b\right\rangle.\end{array}$
@@
$\begin{array}[]{rl}\dot{\mathcal{E}}_{\gamma}(t)&=\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*},\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle\\ &~~~-\frac{t\beta(t)}{\alpha-1}\left\langle\lambda^{*},Ax_{\gamma}(t)-b\right\rangle-\frac{\rho t\beta(t)}{\alpha-1}\left\|Ax_{\gamma}(t)-b\right\|^{2}\\ &\leq\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~-\frac{t\beta(t)}{\alpha-1}\left(f(x_{\gamma}(t))-f(x^{*})\right)-\frac{t\beta(t)}{\alpha-1}\left(g_{\gamma}(x_{\gamma}(t))-g_{\gamma}(x^{*})\right)\\ &~~~-\frac{t\beta(t)}{\alpha-1}\left\langle\lambda^{*},Ax_{\gamma}(t)-b\right\rangle-\frac{\rho t\beta(t)}{\alpha-1}\left\|Ax_{\gamma}(t)-b\right\|^{2}\\ &=\left(\frac{(3-\alpha)t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~-\frac{\rho t\beta(t)}{2(\alpha-1)}\left\|Ax_{\gamma}(t)-b\right\|^{2},\end{array}$ (15)
where the inequality holds due to the convexity of $f$ and $g_{\gamma}$. Since $\sup\limits_{t\geq t_{0}}\frac{t\dot{\beta}(t)}{\beta(t)}\leq\alpha-3$, we have $\frac{(3-\alpha)t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\leq 0$, $\forall t\geq t_{0}$. Then, $\dot{\mathcal{E}}_{\gamma}(t)\leq 0$, $\forall t\geq t_{0}$. This means that $\mathcal{E}_{\gamma}(t)\leq\mathcal{E}_{\gamma}(t_{0})$, $\forall t\geq t_{0}$. For any $t\geq t_{0}$, it follows from integrating (15) from $t_{0}$ to $t$ that
@@
$|(f+g)(x_{\gamma}(t))-(f+g)(x^{*})|$
$\leq\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})+|\langle\lambda^{*},Ax_{\gamma}(t)-b\rangle|+\frac{\rho}{2}\|Ax_{\gamma}(t)-b\|^{2}$
$\leq\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax_{\gamma}(t)-b\|+\frac{\rho}{2}\|Ax_{\gamma}(t)-b\|^{2}.$
Then,
$|(f+g)(x_{\gamma}(t))-(f+g)(x^{*})|=\mathcal{O}\left(\frac{1}{t\sqrt{\beta(t)}}\right),~\textup{as}~t\to+\infty.$
By the properties of the Moreau-Yosida regularization reported in (L1997, , Section 2), there exists a subsequence $\left\{\left(x_{\gamma}(t),{\lambda}_{\gamma}(t)\right)\right\}_{\gamma>0}$ of solution of System (13) that converges to the solution of System (10). Thus, we obtain the desired results by passing to the limit as $\gamma\rightarrow 0$.

### v2

@@ 1 Introduction
$\left\{\begin{split}&\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}~~{f(x)}\\ &\mbox{s.t.}~~Ax=b,\end{split}\right$. (1)
@@
$\left\{\begin{split}&\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)+\theta t\dot{\lambda}(t)\right)=0,\\ &\ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta t\dot{x}(t),\lambda(t)\right)=0,\end{split}\right$. (2)
@@
$\left\{\begin{split}&\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)+\theta t\dot{\lambda}(t)\right)=0,\\ &\ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta t\dot{x}(t),\lambda(t)\right)=0,\end{split}\right$. (3)
@@
$\left\{\begin{split}&\ddot{x}(t)+r\dot{x}(t)+\beta(t)\nabla_{x}\mathcal{L}_{\sigma}\left(x(t),\lambda(t)\right)=0,\\ &\dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\sigma}\left(x(t)+\theta\dot{x}(t),\lambda(t)\right)=0,\end{split}\right$. (4)
@@
$\left\{\begin{split}&\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}~~{f(x)+g(x)}\\ &\mbox{s.t.}~~Ax=b,\end{split}\right$. (5)
@@
$\left\{\begin{split}&\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\partial_{x}\mathcal{L}_{\rho}\left(x(t),\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t)\right)\ni 0,\\ &\ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}_{\rho}\left(x(t)+\frac{t}{\alpha-1}\dot{x}(t),\lambda(t)\right)=0,\end{split}\right$. (6)
@@ 2 Preliminaries
$(x^{*},\lambda^{*})\in\mathcal{S}\Leftrightarrow\left\{\begin{split}&0\in\partial_{x}\mathcal{L}_{\rho}(x^{*},\lambda^{*})=\nabla f(x^{*})+\partial g(x^{*})+A^{*}\lambda^{*},\\ &0=\nabla_{\lambda}\mathcal{L}_{\rho}(x^{*},\lambda^{*})=Ax^{*}-b,\end{split}\right$.
@@ 3 Fast Convergence Rates for Differential System (6)
$\left\{\begin{split}&\ddot{x}_{\gamma}(t)+\frac{\alpha}{t}\dot{x}_{\gamma}(t)+\beta(t)\Big(\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\\ &~~~~+A^{*}\left(\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right)+\rho A^{*}\left(Ax_{\gamma}(t)-b\right)\Big)=0,\\ &\ddot{\lambda}_{\gamma}(t)+\frac{\alpha}{t}\dot{\lambda}_{\gamma}(t)-\beta(t)\left(A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b\right)=0.\end{split}\right$. (9)
@@
$\mathcal{E}_{\gamma}(t)=$ $\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$ (10)
$+\frac{1}{2}\lVert x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\rVert^{2}+\frac{1}{2}\lVert\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\rVert^{2}.$
Clearly, $\mathcal{E}_{\gamma}(t)\geq 0$ for all $t\geq t_{0}$. Then,
$\begin{split}\dot{\mathcal{E}}_{\gamma}(t)=&\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}\left(x_{\gamma}(t),\lambda^{*}\right)-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\langle\nabla_{x}\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*}),\dot{x}_{\gamma}(t)\rangle\\ &+\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle\\ &+\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle\\ =&\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}\left(x_{\gamma}(t),\lambda^{*}\right)-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right),\dot{x}_{\gamma}(t)\right\rangle+\frac{t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle\lambda^{*},A\dot{x}_{\gamma}(t)\right\rangle\\ &+\frac{\rho t^{2}\beta(t)}{(\alpha-1)^{2}}\left\langle Ax_{\gamma}(t)-b,A\dot{x}_{\gamma}(t)\right\rangle\\ &+\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle\\ &+\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle.\end{split}$}
@@
$\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{x}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{x}_{\gamma}(t)\right\rangle$
$=$ $-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle$
$-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),A^{*}\left(\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right)+\rho A^{*}\left(Ax_{\gamma}(t)-b\right)\right\rangle$
$=$ $-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*}+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t),\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle$
$-\frac{t\beta(t)}{\alpha-1}\left\langle A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b,\lambda_{\gamma}(t)+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t)\right\rangle$
$-\frac{\rho t\beta(t)}{\alpha-1}\left\langle A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b,Ax_{\gamma}(t)-b\right\rangle,$
@@
$\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),\frac{\alpha}{\alpha-1}\dot{\lambda}_{\gamma}(t)+\frac{t}{\alpha-1}\ddot{\lambda}_{\gamma}(t)\right\rangle$
$=$ $\frac{t\beta(t)}{\alpha-1}\left\langle\lambda_{\gamma}(t)-\lambda^{*}+\frac{t}{\alpha-1}\dot{\lambda}_{\gamma}(t),A\left(x_{\gamma}(t)+\frac{t}{\alpha-1}\dot{x}_{\gamma}(t)\right)-b\right\rangle.$
@@
$\dot{\mathcal{E}}_{\gamma}(t)=$ $\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$ (11)
$-\frac{t\beta(t)}{\alpha-1}\left\langle x_{\gamma}(t)-x^{*},\nabla f\left(x_{\gamma}(t)\right)+\nabla g_{\gamma}\left(x_{\gamma}(t)\right)\right\rangle$
$-\frac{t\beta(t)}{\alpha-1}\left\langle\lambda^{*},Ax_{\gamma}(t)-b\right\rangle-\frac{\rho t\beta(t)}{\alpha-1}\left\|Ax_{\gamma}(t)-b\right\|^{2}$
$\leq$ $\left(\frac{2t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$
$-\frac{t\beta(t)}{\alpha-1}\left(f(x_{\gamma}(t))-f(x^{*})\right)-\frac{t\beta(t)}{\alpha-1}\left(g_{\gamma}(x_{\gamma}(t))-g_{\gamma}(x^{*})\right)$
$-\frac{t\beta(t)}{\alpha-1}\left\langle\lambda^{*},Ax_{\gamma}(t)-b\right\rangle-\frac{\rho t\beta(t)}{\alpha-1}\left\|Ax_{\gamma}(t)-b\right\|^{2}$
$=$ $\left(\frac{(3-\alpha)t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\right)\left(\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$
$-\frac{\rho t\beta(t)}{2(\alpha-1)}\left\|Ax_{\gamma}(t)-b\right\|^{2},$
where the inequality holds due to the convexity of $f$ and $g_{\gamma}$. Since $\sup\limits_{t\geq t_{0}}\frac{t\dot{\beta}(t)}{\beta(t)}\leq\alpha-3$, we have $\frac{(3-\alpha)t\beta(t)}{(\alpha-1)^{2}}+\frac{t^{2}\dot{\beta}(t)}{(\alpha-1)^{2}}\leq 0$ for all $t\geq t_{0}$. Then, $\dot{\mathcal{E}}_{\gamma}(t)\leq 0$ for all $t\geq t_{0}$. This means that $\mathcal{E}_{\gamma}(t)\leq\mathcal{E}_{\gamma}(t_{0})$ for all $t\geq t_{0}$. For any $t\geq t_{0}$, it follows from integrating (11) from $t_{0}$ to $t$ that
@@
$\begin{split}&|(f+g)(x_{\gamma}(t))-(f+g)(x^{*})|\\ \leq&\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})+|\langle\lambda^{*},Ax_{\gamma}(t)-b\rangle|+\frac{\rho}{2}\|Ax_{\gamma}(t)-b\|^{2}\\ \leq&\mathcal{L}_{\rho}(x_{\gamma}(t),\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax_{\gamma}(t)-b\|+\frac{\rho}{2}\|Ax_{\gamma}(t)-b\|^{2}.\end{split}$
Then, $|(f+g)(x_{\gamma}(t))-(f+g)(x^{*})|=\mathcal{O}\left(\frac{1}{t\sqrt{\beta(t)}}\right),~\textup{as}~t\to+\infty.$ By the properties of the Moreau-Yosida regularization reported in (L1997, , Section 2), there exists a subsequence $\left\{\left(x_{\gamma}(t),{\lambda}_{\gamma}(t)\right)\right\}_{\gamma>0}$ of solution of System (9) that converges to the solution of System (6). Thus, we obtain the desired results by passing to the limit as $\gamma\rightarrow 0$.